Say me about motion of rigid bodies and derivation of moment of inertias of different types of objects like cylinder, rod,e.t.c

Grade:11

1 Answers

Vikas TU
7 years ago
Moment of inertia for the rigid body is simply as: => product of their masses with the square of the distance from point of interest.
For cyclinder,
dI = dM*r^2 where r is the radii of the cylinder 
dM = p*dV and dV = dA * L and dA = 2pirdr and thus dV = 2pirLdr
substituting all the values and integerating we get the M.O.I for the cylinder as=> MR^2/2 for solid
and MR^2 for Hollow cylinder respetively.